OpenStudy (anonymous):

Solve for x: 1/3x - 3 = 2x + 4

OpenStudy (texaschic101):

1/3x - 3 = 2x + 4 --- multiply the entire equation by 3 x - 9 = 6x + 12 -- subtract 6x from both sides x - 6x - 9 = 12 -- add 9 to both sides x - 6x = 12 + 9 -- combine like terms -5x = 21 x = -21/5

OpenStudy (anonymous):

can you work it out please

OpenStudy (nirmalnema):

is x in numenator of denominator..in 1/3x

OpenStudy (texaschic101):

I thought it was (1/3) x.....if it is not, then my answer is wrong